OpenStudy (anonymous):

Find the derivative of y ith espect to x: y=(2x+1) ln(x-1)

OpenStudy (anonymous):

Use the product rule. u=2x+1, v=ln(x-1) \[d(u*v)/dx \ = \ u*v' + v*u' = (2x+1) * 1/(x-1) + \ln(x-1) * 2\] and you can simplify after that.

OpenStudy (amistre64):

u' = 2 and v' = 1/(x-1) if I recall correctly...